1. A group of 463 first-year college students were asked, “About how many hours do you study during a typical week?” The mean response is 15.3 hours. Assume that the study time is normally distributed with a population standard deviation of 8.5 hours. Construct a 99% confidence level interval for the mean study time of all first-year students.
Mean Proportion

Answer by Boreal(15235)

99%CI width is z(0.995)*sd/sqrt(n)=2.576*8.5/sqrt(463)=1.02
the interval is (7.48,9.52)